# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ 4.2.3 Multi-Object and Scene Generation
Multi-Object Generation: Several studies [98, 78] have explored the generation of multiple composite objects, which not only concentrate on the individual objects but also aim to investigate the interactions between multiple objects.
CG3D [98] discusses both aspects separately. For the reconstruction of individual objects, CG3D introduces the $K$ nearest neighbors loss on Alpha hull [99] to ensure that the predicted Gaussian primitives are uniformly distributed and concentrated on the object surfaces. For predicting the interactions between multiple objects, CG3D leverages SDS and probabilistic graph models extracted from text to predict the relative relationships between objects. Finally, by incorporating priors such as gravity and contact relationships between objects, CG3D achieves models with realistic physical interactions.
To simplify such a problem, GALA3D [78] uses layouts generated by large language models (LLMs) to guide multiple object reconstructions. By exploring the distribution of positions and optimizing the shapes of Gaussian primitives according to the layout, GALA3D generates scenes that conform to the specified layouts. Additionally, by supervising the generation of individual objects and entire scenes through SDS and introducing a Layout Refinement module, GALA3D achieves more realistic and text-consistent generation results.
Scene Generation: Unlike object-centric generation, scene generation often requires the incorporation of additional auxiliary information, such as pre-trained monocular depth estimation models, to achieve high-precision initialization.
To achieve this, LucidDreamer2 [100] designs a two-stage generative paradigm. In the first stage, LucidDreamer2 leverages pre-trained text-to-image models and monocular depth estimation models to initialize the point clouds, and introduces Stable Diffusion inpainting models [101] to complete a multi-view consistent scene point cloud. In the second stage, the generated point cloud is used to initialize 3DGS, with extended supervision images to ensure a smoother training process. Based on a similar paradigm, Text2Immersion [102] introduces a poses-progressive generation strategy to achieve a more stable training process, and incorporates enlarged viewpoints and pre-trained super-resolution models to optimize the generated scenes.
4.2.4 4D Generation
In addition to static scenes, some studies [103, 76, 104] have begun to delve deeply into dynamic 3D scenes. Analogous to static scene generation using text-to-image SDS, it is natural to consider that text-to-video SDS could potentially generate dynamic scenes.
Align Your Gaussians (AYG) [103] explicitly divides the problem into two stages: static 3DGS reconstruction and 4DGS (dynamic 3DGS) reconstruction. In the static reconstruction stage, AYG incorporates the pre-trained text-guided multiview diffusion model MVDream [105] and text-to-image model to jointly supervise the 3DGS training. In the dynamic reconstruction stage, AYG proposes using both pre-trained text-to-image and text-to-video models to supervise the dynamic 3DGS training. Moreover, AYG introduces a simplified score distillation loss to reduce training uncertainty.
Building on DreamGaussian [75], DreamGaussian4D [76] takes a single reference image as input and utilizes pre-trained image-to-video models as well as multi-view generation models for supervised training.
GaussianFlow [104] aims to use optical flow information from videos as auxiliary supervision to assist 4DGS creation. Similar to the dynamic Gaussian work, Motion4D [106], this study first analyzes the relationship between Gaussian primitives’ movement in 3D space and pixel movement (optical flow) in 2D pixel space. By aligning the motion of 3D Gaussians with optical flow, GaussianFlow can achieve text-to-4DGS and image-to-4DGS generation.
However, the instability of video generation models affects the generation performance in the SDS-based paradigm.
4DGen [107] addresses this issue by introducing multi-view generation models for each frame of the given video to create pseudo-labels, replacing video generation models. To ensure temporal consistency, 4DGen utilizes multi-scale intermediate representations from Hexplane [108] to constrain the smoothness of Gaussian primitives over time, further enhancing 4DGS generation quality. Similarly, in the same context and experimental setup, Fast4D [109] forms a matrix of images using these pseudo-labels and considers continuity in both temporal and spatial dimensions (i.e., rows and columns). Using the image matrix as supervision, the Fast4D propose a novel time-dependent 3DGS representation [110] to achieve efficient and high-quality generation.
Additionally, some studies [111] focus on animating existing static 3DGS.
@@
For animating existing 3DGS from input video, BAGS [111] introduces neural bones and skinning weights to describe the spatial deformation based on canonical space. Using diffusion model priors and rigid body constraints, BAGS can be manually manipulated to achieve novel pose rendering.
4.3 Autonomous Driving
In the field of autonomous driving, 3DGS is primarily applied to the dynamic reconstruction of large-scale driving scenes and combined SLAM applications.
4.3.1 Autonomous Driving Scene Reconstruction
Reconstruction driving scenes is a challenging task, involving multiple technical domains such as large-scale scene reconstruction, dynamic object reconstruction, static object reconstruction, and Gaussian mixture reconstruction.
A substantial body of works [112, 113, 114] categorizes the reconstruction process into static background reconstruction and dynamic target reconstruction.
DrivingGaussian [112] aims to utilize multi-sensor data for reconstructing large-scale dynamic scenes in autonomous driving. In the static background, DrivingGaussian introduces incremental static 3D Gaussians under different depth bins to mitigate the scale confusion caused by distant street scenes. For dynamic objects, DrivingGaussian introduces dynamic Gaussian graphs to construct relationships among multiple targets (whose attributes include position, local-to-world coordinate transformation matrices, and orientation, etc.), jointly reconstructing the entire autonomous driving scene with the static background. StreetGaussians [113] adopts a similar approach, with the key difference being the introduction of semantic attributes during the reconstruction of background and foreground. Also, StreetGaussians uses Fourier transforms to efficiently represent the SH’s temporal changes of dynamic 3DGS. Building on previous works, HUGS [114] incorporates the Unicycle Model and the modeling of forward and angular velocities to assist in the dynamic reconstruction under physical constraints. Similar to previous dynamic 3DGS efforts [115, 106], HUGS also employs optical flow supervision, combined with rendered RGB loss, semantic loss, and Unicycle Model losses, thus improving dynamic reconstruction accuracy.
Moreover, 3DGS have been applied to multimodal spatiotemporal calibration tasks [116].
@@ 4.3.1 Autonomous Driving Scene Reconstruction
By leveraging the LiDAR point cloud as a reference for the Gaussians’ positions, 3DGS-Calib [116] constructs a continuous scene representation and enforces both geometrical and photometric consistency across all sensors, achieving accurate and robust calibration with significantly reduced training times compared to NeRF-based methods.
4.3.2 Simultaneous Localization and Mapping (SLAM)
SLAM is a fundamental problem in robotics and computer vision, where a device constructs a map of an unknown environment while simultaneously determining its own location within that environment. The technical approaches to SLAM can be broadly categorized into traditional methods, techniques involving NeRF, and approaches related to 3DGS. Among these, 3DGS methods stand out for their ability to offer continuous surface modeling, reduced memory requirements, improved noise and outlier handling, enhanced hole filling and scene repair, and flexible resolution in 3D mesh reconstruction [117].
Some studies [118, 119, 120, 121, 122] have retained the traditional SLAM inputs and approached this from two perspectives: online tracking and incremental mapping.
In earlier works, GS-SLAM [118] utilizes 3DGS as a scene representation for SLAM and introduces an adaptive expansion strategy. This strategy involves the dynamic addition of new Gaussian primitives during the training phase and the removal of noisy primitives based on captured depth and rendered opacity, thereby facilitating continuous scene reconstruction during motion. For camera tracking, GS-SLAM proposes an advanced coarse-to-fine optimizing strategy. Initially, a sparse set of pixels is rendered to optimize the tracking loss and obtain an initial, coarse estimate of the camera poses. Subsequently, based on these coarse camera poses and depth observations, reliable Gaussian primitives are selected in 3D space to guide GS-SLAM in re-rendering regions with well-defined geometric structures, which further refines the coarse camera poses. The entire process is supervised using a re-rendering loss.
Photo-SLAM [119] introduces a novel SLAM framework featuring a Hyper Primitives Map, which incorporates ORB features [123] along with Gaussian attributes. Based on such representation, this framework leverages the Levenberg-Marquardt (LM) algorithm [124] to optimize both the localization and geometric maps from projection relationships. Building on these results, Photo-SLAM proposes a geometry-based densification strategy and a Gaussian-Pyramid-Based Learning mechanism to construct photorealistic mapping. Finally, the framework integrates loop closure [123] to further correct camera poses and improve mapping quality.
Gaussian-SLAM [121] addresses the mapping challenge by dividing the map into multiple submaps, each reconstructed separately, thereby alleviating catastrophic forgetting. For camera tracking, the authors observe that tracking accuracy is constrained by the extrapolation capabilities of 3DGS and recommend incorporating trajectory assistance from DROID-SLAM [125] to enhance the reconstruction.
The work [126] analytically derives the Jacobian matrix concerning camera poses, proposing an efficient camera pose optimization strategy. During the mapping phase, the authors estimate frame-to-frame co-visibility and design a keyframe selection and management mechanism that prioritizes frames with low co-visibility as keyframes. This ensures the efficient use of non-redundant keyframes in the same region, thereby enhancing mapping efficiency.
Based on previous works, RGBD GS-ICP SLAM [127] integrates the Generalized Iterative Closest Point (G-ICP) [128] algorithm into the mapping and tracking. By sharing the covariances between G-ICP and 3DGS as well as employing scale alignment techniques, this work minimizes redundant computations and facilitates rapid convergence.
Similarly, the work by Sun et al. [129] also proposes a densification strategy guided by holes and rendering errors to map unobserved areas and refine re-observed regions. Additionally, the authors design a new regularization term to alleviate the problem of catastrophic forgetting.
Semantics are important as they not only provide scene understanding but also stabilize the training process.
SGS-SLAM [122] employs multi-channel geometric, appearance, and semantic features for rendering and optimization and proposes a keyframe selection strategy based on geometric and semantic constraints to enhance performance and efficiency. Furthermore, due to the construction of semantic representations, SGS-SLAM is capable of editing 3DGS, which correspond to semantically consistent regions. Based on it, SEMGAUSS-SLAM [130] designs a feature-level supervision for robustness and introduces a feature-based Bundle Adjustment to mitigate cumulative drift during tracking. Subsequent works, NEDS-SLAM [131], also adopts this concept, introducing semantic feature-assisted SLAM optimization and incorporating DepthAnything [132] to learn semantically rich features with 3D spatial awareness. Additionally, NEDS-SLAM proposes a novel pruning method based on virtual multi-view consistency checks to identify and eliminate outliers.
Additionally, there are several works focusing on related issues such as localization [133] and navigation [134].
3DGS-ReLoc [133] introduces a LiDAR data initialization to assist in 3DGS mapping. Building upon this, it proposes dividing the 3DGS submap into a 2D voxel map and utilizing a KD-tree for efficient spatial queries, thereby mitigating excessive GPU memory consumption. Finally, 3DGS-ReLoc employs a feature-based matching and Perspective-n-Point (PnP) method to iteratively refine the pose of the query image, achieving precise localization within the global map.
@@ 4.3.2 Simultaneous Localization and Mapping (SLAM)
In the context of indoor navigation, GaussNav [134] focuses on the instance image navigation (IIN) task. Based on reconstructed 3DGS maps, GaussNav proposes an image target navigation algorithm, achieving impressive performance through classification, matching, and path planning.
5 Extensions of 3D Gaussian Splatting
As an essential technique for 3D representation, 3DGS can be further expanded for more capabilities, including dynamic 3DGS [Fig. 5(a)], surface representation from 3DGS [Fig. 5(c)], editable 3DGS [Fig. 5(b)], 3DGS with semantic understanding [Fig. 5(d)], and 3DGS-based physics simulation [Fig. 5(e)].
(a) Dynamic 3DGS [135].
(b) Editable 3DGS [136].
(c) Surface Representation [137].
(d) Semantic Understand [138].
(e) Physics Simulation [139].
Figure 5: Five different Extensions of 3DGC.
5.1 Dynamic 3D Gaussian Splatting
The study of dynamic 3DGS has recently garnered significant attention from researchers. The reconstruction of dynamic scenes transcends the limitations of static scene reconstruction and can be effectively applied to fields such as human motion capture and autonomous driving simulation. Unlike static 3DGS, dynamic 3DGS must account for consistency not only in the spatial dimension but also in the temporal dimension, ensuring continuity and smoothness over time. Here, we categorize them into Multi-view Videos and Monocular Video based on different reconstruction inputs.
5.1.1 Multi-view Videos
Some works [135, 140] attempted to directly construct dynamic 3DGS frame by frame.
An early work [135] extends 3DGS from static to dynamic scenes by allowing the Gaussians to move and rotate over time while maintaining persistent attributes such as color, opacity, and size. The reconstruction is performed temporally online, where each timestep is initialized using the previous timestep’s representation. The first timestep acts as an initialization for optimizing all attributes, which are then fixed for subsequent timesteps except for those defining motion. Physical priors, including local rigidity, local rotational-similarity, and long-term local-isometry, regularize the motion and rotation of the Gaussians, as shown in Eq. 5-7. The authors also discuss the impact of factors such as similar background information and multi-view camera differences on reconstruction performance.
Similarly, 3DGStream [140] designs a two-stage training process for transformation predictions. In the first stage, it introduces Neural Transformation Cache and I-NGP [9] to reconstruct dynamic 3DGS. In the second stage, 3DGStream proposes an adaptive densification strategy, initializing new Gaussian positions by calculating gradients.
Other works [141, 142] aim to achieve such performance by predicting deformations.
SWAGS [141] introduces a window-based 4DGS, sampling videos into multiple windows to achieve long-term scene reconstruction. To ensure that the deformation degree within each window is as similar as possible, an adaptive window division method based on average flow is introduced. A dynamic MLP is then used to guide the optimization to focus on dynamic areas. On the basis of predictions from multiple sampled windows, SWAGS proposes using the consistency of overlapping frames from adjacent windows to design a self-supervised loss that fine-tunes the entire scene, thus eliminating temporal discontinuities from window divisions.
5.1.2 Monocular Video
Some monocular dynamic 3DGS works [143, 144, 145, 146, 106] tend to divide into two stages: canonical reconstruction and deformation prediction.
The study [143] first reconstructs static 3DGS in the canonical space. Then, it takes the encoded positions and time $t$ as inputs, outputting offsets in terms of positions, rotations, and scales. To mitigate overfitting issues caused by inaccurate poses, this work introduces an annealing smooth training paradigm, which incorporates linearly decaying Gaussian noise.
Building on this foundation, GauFRe [144] proposes a paradigm that decouples dynamic and static scene modeling, where the dynamic part utilizes a method similar to that of [143]. 4D-GS [145] introduces the multi-scale HexPlane [108] as the foundational representation to encode temporal and spatial information. To optimize the training process, 4D-GS employs multi-head decoders to predict different attributes of Gaussian primitives separately. MD-Splatting [146] also builds on this by incorporating the local rigidity loss and isometric loss as proposed in [135], and designs a regularization term based on the law of conservation of momentum, making the dynamic motion trajectories smoother. Additionally, the authors include shadow prediction in the decoding part, further enhancing the realism of the reconstruction.
The work by Guo et al. [106] constructs a flow-augmentation method by analyzing the correspondence between 3D Gaussian motion and pixel-level flow, introducing additional optical flow-based supervision based on uncertainty and dynamic awareness. Additionally, it proposes a Motion Injector and Dynamic Map Refinement strategy based on a velocity field to mitigate the challenges associated with predicting deformations.
Instead of discrete offsets, exploring temporally continuous motion can promote smoothness in the time dimension.
The work by Katsumata et al. [115] incorporates Fourier approximations over time and optical flow supervision to model the attribute changes of 3DGS, ensuring continuous changes without introducing excessive parameters.
DynMF [147] argues that every scene is comprised of a limited fixed number of trajectories and introduces efficient basis functions for spatial motion. Based on these basis functions, DynMF designs motion coefficients to predict dynamic positions and rotation attributes, and introduces constraints such as sparsity and rigidity to further enhance optimization performance. Similarly, STG [148] proposes the use of temporally relevant Radial Basis Functions to represent the opacity attributes, and utilizes parameterized polynomials to describe the motion and rotation of Gaussian primitives. Additionally, a new time- and view-dependent feature splatting paradigm is introduced to replace spherical harmonics for color.
@@ 5.1.2 Monocular Video
Gaussian-Flow [149] aims to develop a representation capable of fitting variable motion by analyzing the advantages and disadvantages of polynomial [147, 148] and Fourier series fitting [115]. It then proposes a model that combines both methods to construct a dual-domain deformation prediction model. Finally, Gaussian-Flow introduces an adaptive time-step scaling strategy and a temporal smoothing and rigid loss to enhance training stability and temporal continuity.
SC-GS [150] aims to compress motion representations in dynamic scenes based on sparse control points. By predicting the positions of control points, radial-basis-function (RBF) kernels, and transformations, the motion of the entire dynamic field over time is calculated using the Linear Blend Skinning [151]. Thanks to this strategy, SC-GS proposes an adaptive pruning and cloning strategy for control points based on neighboring point RBF weights and gradients, and it demonstrates strong applicability in 3DGS editing.
Recent works [110, 152] aim to extend 3DGS to 4D space for dynamic 3D scenes representation.
The work [110] achieves end-to-end training by incorporating the time dimension into the 3D representation directly. The core idea is to simultaneously consider spatial variables ($XYZ$) and time variable ($T$) along with their interrelations, rather than treating them as independent variables. For 4D Gaussian primitives, it designs a 4D representation of the rotation matrix $\bm{R}\in\mathbb{R}^{4\times 4}$ and scaling factor $\bm{S}$ (4D diagonal), and employs 4D spherical harmonics to further describe the color variations over time.
Similarly, the work [152] proposes a rotor-based 4D Gaussian Splatting (4DGS) representation, wherein the rotation attributes of 4DGS are expressed by decomposing 4D rotors into eight components. These components, along with corresponding parameters, are used to describe the rotation in space-time. When representing dynamic 3DGS, it slices 4DGS at different timestamps, effectively addressing the sudden appearance or disappearance of objects in highly dynamic scenes. Furthermore, the approach enforces consistency in 4D space through the introduction of a 4D Consistency Loss.
5.2 Surface Representation
Although 3DGS enables highly realistic rendering, extracting surface representations remains challenging. After the optimization process, the resulting representations often lack ordered structures and do not correspond well to the actual surfaces. However, mesh-based representations remain the preferred choice in many workflows, as they allow the use of robust tools for editing, sculpting, animation, and relighting.
When it comes to surface reconstruction, Signed Distance Functions (SDF) are an indispensable topic.
NeuSG [153] attempts to jointly optimize NeuS [154] and 3DGS and introduces several regularization terms, including Scale Regularization, Normal Regularization, and Eikonal Regularization [155], to ensure that the 3DGS are as flat as possible and distributed along the target surface.
SuGaR [156] leverages the 3DGS surface characteristics to design an idealized SDF. This idealized representation is then used to constrain the actual predicted SDF and its normals, thereby encouraging the optimized Gaussians to more closely align with the object’s surface. Then, using Poisson reconstruction, SuGaR extracts a mesh from the aligned 3DGS, which is faster and more scalable than the Marching Cubes algorithm [157]. Additionally, an optional refinement step binds new Gaussians to mesh and jointly optimizes them, enabling high-quality rendering and surfaces.
And then, 3DGSR [158] aims to integrate neural implicit SDF with 3DGS by designing a differentiable SDF-to-opacity transformation function, enabling the optimization of 3DGS to update the SDF. To address the challenge of optimizing continuous SDF with discrete 3DGS, 3DGSR proposes enforcing consistency between the depth (normals) obtained from volumetric rendering and those inferred from 3DGS.
GSDF [159] similarly introduces consistency constraints and employs a dual-stream network combining 3DGS and neural implicit SDF based on Scaffold-GS [26] and NeuS [154]. To improve the sampling efficiency of SDF, GSDF uses the depth maps from the 3DGS branch to guide the ray sampling process, and to improve the distribution of Gaussian primitives on surfaces, designs a Geometry-aware Gaussian Density Control for 3DGS based on the SDF branch.
Other studies [160, 161, 137] aim to address this issue by enhancing the intrinsic attributes of 3DGS.
The work [160] introduces a novel representation termed Gaussian Surfels, which exhibits an enhanced capability for surface reconstruction. Based on this representation, it proposes a depth-normal consistency loss to address the gradient vanishing problem and a volumetric cutting strategy to prune unnecessary voxels in regions with depth errors and discontinuities. Finally, this work applies screened Poisson reconstruction to generate the surface mesh.
Gaussian Opacity Fields (GOF) [161] are developed based on 3DGS, wherein 3DGS is normalized along the ray to form a 1DGS for volume rendering. GOF also incorporates depth distortion and normal consistency losses, facilitating the extraction of surface meshes from tetrahedral grids.
Similarly, 2D Gaussian Splatting [137] (2DGS) replaces 3DGS with planar disks to represent surfaces, which are defined within the local tangent plane. In the rendering process, 2DGS abandons direct affine transformations and uses three non-parallel planes to define the ray-splat intersections, followed by rasterization after applying a low-pass filter.
5.3 Editable 3D Gaussian Splatting

### v2

@@ IV-B3 Multi-Object and Scene Generation
Multi-Object Generation: Several studies [126, 98, 127, 128] have explored the generation of multiple composite objects, which not only concentrate on the individual objects but also aim to investigate the interactions between multiple objects. For predicting the interactions between multiple objects, CG3D [126] leverages SDS and probabilistic graph models extracted from text to predict the relative relationships between objects and incorporates priors such as gravity and contact relationships between objects, CG3D achieves models with realistic physical interactions.
Scene Generation: Unlike object-centric generation, scene generation typically requires the incorporation of additional information, such as pre-trained depth estimation models [129, 130] or Large Language Models [131, 127], where the former provides spatial understanding for projecting images into 3D space, while the latter enhances the quality of text prompts. LucidDreamer2[129] employs a two-stage approach: first initializing point clouds using text-to-image and depth estimation models with inpainting[132] for consistency, then converting to 3DGS with extended image supervision.
IV-B4 4D Generation
Analogous to static scene generation using text-to-image SDS, it is natural to consider that text-to-video SDS could potentially generate dynamic scenes [133, 96, 134, 135, 136, 137]. These works primarily focus on designing video-based SDS losses or exploring hybrid supervision with T2I (Text to Image) and T2V (Text to Video) models. As an example, Align Your Gaussians [133] adopts a two-stage approach: first reconstructing static 3DGS using MVDream[138] and text-to-image supervision, then extending to 4DGS with text-to-video guidance and simplified score distillation loss. Although these methods are effective, the inherent limitations of the aforementioned SDS-based paradigm persist.
To mitigate this issue, subsequent works focus on generating pseudo-labeled images from additional views to facilitate dynamic 3DGS reconstruction [139, 140, 141, 142]. Among them, 4DGen [139] generates multi-view pseudo-labels per frame, while employing Hexplane’s [143] multi-scale features to maintain temporal consistency in 4DGS generation.
@@
Furthermore, some studies focus on animating static canonical 3DGS [144, 145] to achieve better control. Among them, BAGS [144] introduces neural bones and skinning weights to describe the spatial deformation based on canonical space. Using diffusion model priors and rigid body constraints, BAGS can be manually manipulated to achieve novel pose rendering.
IV-C Autonomous Driving
IV-C1 Autonomous Driving Scene Reconstruction
Reconstruction driving scenes is a challenging task, involving multiple technical domains such as large-scale scene reconstruction, dynamic object reconstruction, static object reconstruction, and Gaussian mixture reconstruction. Existing works [146, 147, 148] partitions the entire process into static background and dynamic target reconstruction. DrivingGaussian [146] reconstructs large-scale driving scenes by combining depth-binned static 3DGS for background and dynamic Gaussian graphs for multiple targets, utilizing multi-sensor data. Then StreetGaussians [147] extends this approach by incorporating semantic attributes and employing Fourier transforms for efficient SH temporal modeling in dynamic 3DGS. Subsequent works [149], aiming to further improve reconstruction efficiency, attempt to directly reconstruct entire scenes by Tightly Coupled LiDAR-Camera Gaussian Splatting.
@@ IV-C1 Autonomous Driving Scene Reconstruction
Moreover, 3DGS have been applied to multimodal spatiotemporal calibration tasks [150]. By leveraging the LiDAR point cloud as a reference for the Gaussians’ positions, 3DGS-Calib [150] constructs a continuous scene representation and enforces both geometrical and photometric consistency across all sensors, achieving accurate and robust calibration with improved efficiency compared to NeRF-based works.
IV-C2 Simultaneous Localization and Mapping (SLAM)
SLAM is a key problem in robotics and computer vision, where a device builds a map of an unknown environment while locating itself within it. Some studies [151, 152, 153, 154, 155, 156] have retained the traditional inputs and approached this from two perspectives: online tracking and incremental mapping. In early work, GS-SLAM [151] utilizes 3DGS for SLAM with adaptive primitive expansion and employs a coarse-to-fine optimization strategy: first optimizing camera poses using sparse pixels, then refining them through selective re-rendering of reliable Gaussians. In parallel, Photo-SLAM [152] combines ORB features [157] and Gaussian attributes in a Hyper Primitives Map, utilizing LM optimization [158] and loop closure [157] for photorealistic SLAM reconstruction. While these approaches achieve higher efficiency than NeRF-based methods, further optimization of computational performance is essential for real-world deployment. Therefore, CG-SLAM [155] leverages an uncertainty-aware 3D Gaussian field and a GPU-accelerated framework, achieving superior efficiency. Then, RGBD GS-ICP SLAM [156] enhances efficiency by integrating G-ICP [159] with shared covariances and scale alignment techniques for faster convergence. However, these methods remain susceptible to sensor noise in practical applications. Some quantitative reconstruction results are reported in Table III.
Incorporating scene understanding capabilities is equally crucial in SLAM tasks [160, 161, 162], thus prompting several works to integrate semantic information. Among them, SGS-SLAM [160] employs multi-channel geometric, appearance, and semantic features for rendering and optimization and proposes a keyframe selection strategy based on geometric and semantic constraints to enhance performance and efficiency.
@@ IV-C2 Simultaneous Localization and Mapping (SLAM)
Additionally, there are several works focusing on related issues such as localization [163] and navigation [164]. Specifically, 3DGS-ReLoc [163] leverages LiDAR initialization and 2D voxelized submaps with KD-tree for efficient memory usage, while achieving precise localization through feature-based PnP optimization. In the context of indoor navigation, GaussNav [164] focuses on the instance image navigation task. Based on reconstructed 3DGS maps, GaussNav proposes an image target navigation algorithm, achieving impressive performance through classification, matching, and path planning.
TABLE III: Comparison of 3DGS-SLAM on Replica [165].
Method PSNR↑ SSIM↑ LPIPS↓ Tracking RMSE↓
GS-SLAM [166] 34.27 0.98 0.08 0.50
Photo-SLAM [152] 34.96 0.94 0.06 0.60
SplaTAM [153] 34.11 0.97 0.10 0.36
GS-ICP SLAM [156] 38.83 0.98 0.04 0.16
MotionGS [167] 39.60 0.98 0.04 0.49
LoopSplat [168] 36.63 0.99 0.11 0.26
V Extensions of 3D Gaussian Splatting
V-A Dynamic 3D Gaussian Splatting
The study of dynamic 3DGS has recently garnered significant attention. Here, we categorize them into Multi-view Videos and Monocular Video based on different inputs.
V-A1 Multi-view Videos
Some works [169, 170] attempted to directly construct dynamic 3DGS frame by frame. An early work [169] extends 3DGS to dynamic scenes by enabling temporal Gaussian motion while preserving static attributes. It employs online temporal reconstruction with previous-frame initialization and incorporates physical priors, including local rigidity, local rotational-similarity, and long-term local-isometry, for motion regularization, as shown in Eq.4-6. Similarly, 3DGStream [170] employs a two-stage training with Neural Transformation Cache and I-NGP [171] for dynamic reconstruction, followed by gradient-based adaptive Gaussian densification. Despite promising results, these methods are limited to reconstructing only scene elements visible in the initial frame. To address this limitation, other works [172, 173] aim to achieve such performance by predicting deformations. SWAGS [172] proposes a window-based 4DGS with flow-guided adaptive window division and dynamic MLP optimization, employing inter-window consistency loss for seamless temporal reconstruction.
V-A2 Monocular Video or Multi-view Images
Some works [174, 175, 176, 177, 178, 179, 180] tend to divide into two stages: canonical reconstruction and deformation prediction. The studies [174, 175] reconstruct static 3DGS and predicts temporal deformation in terms of positions, rotations and scales through position-time encoding. Similarly, 4D-GS [176] introduces the multi-scale HexPlane [143] as the foundational representation to encode temporal and spatial information. To further decouple motion and shape parameters in 4D-GS, ST-4DGS [179] introduces a spatial-temporally consistent 4DGS framework that incorporates motion-aware shape regularization and spatial-temporal density control to learn compact 4D representations. Although these methods achieve stable performance, they struggle to handle abrupt motions and sudden object appearances.
@@ V-A2 Monocular Video or Multi-view Images
Instead of discrete offsets, exploring temporally continuous motion can promote smoothness in the time dimension [181, 182, 183, 184]. Gaussian-Flow [183] aims to develop a representation capable of fitting variable motion by analyzing the advantages and disadvantages of polynomial [182, 185] and Fourier series fitting [181]. It then proposes a dual-domain deformation model with adaptive time-step scaling and temporal-rigid constraints for stable and continuous motion prediction.
TABLE IV: Comparison of Dynamic 3DGS on D-NeRF [186].
Method PSNR↑ SSIM↑ LPIPS↓ Train↓ FPS↑
K-Planes [NeRF-based] 31.07 0.97 0.02 54 min 1.20
Deformable3DGS [174] 39.31 0.99 0.01 26 min 85.45
CoGS [187] 37.90 0.983 0.027 - -
SC-GS [184] 43.31 0.997 0.0063 - -
GauFRe [175] 34.5 0.98 0.02 13mins 112
Deformable4DGS [176] 32.99 0.97 0.05 13 min 104.00
RealTime4DGS [188] 32.71 0.97 0.03 10 min 289.07
4DRotorGS [189] 34.26 0.97 0.03 5 min 1257.63
Recent works aim to extend 3DGS to 4D space for dynamic 3D scenes representation. Among them, the work [188] achieves end-to-end 4D training by jointly modeling spatial-temporal variables with 4D Gaussian primitives, incorporating 4D rotation, scaling, and temporal-aware spherical harmonics for color variation. Similarly, the work [189] introduces a rotor-based 4DGS representation with eight-component rotation decomposition, enabling temporal slicing for dynamic objects and enforcing 4D consistency through a dedicated loss. Although these approaches demonstrate robustness in complex scene reconstruction, the compressibility of their representations remains a notable consideration. Quantitative reconstruction results are reported in Table IV.
Dynamic 3DGS reconstruction demonstrate broad applicability across diverse domains, extending beyond previously mentioned human reconstruction IV-A, AIGC IV-B, and autonomous driving IV-C to include robotic manipulation [190], physical simulation [191], and various downstream tasks.
V-B Surface Representation
Although 3DGS enables highly realistic rendering, extracting surface representations remains challenging. In this line of works, Signed Distance Functions are an indispensable topic [192, 193, 194, 195].In early work, SuGaR [192] proposes an idealized SDF to constrain Gaussian surface alignment, enabling efficient mesh extraction through Poisson reconstruction and optional mesh-guided Gaussian refinement for high-quality results. Similarly, 3DGSR [194] integrates neural implicit SDF with 3DGS through a differentiable SDF-to-opacity transformation, maintaining consistency between volumetric and 3DGS-derived depth properties. Another line of research [193, 195] focuses on jointly optimizing NeuS [196] and 3DGS for surfaces. However, these methods exhibit limitations in handling unbounded scenes and computational overhead.
Other studies [197, 198, 199, 200] aim to address this issue by improving 3DGS representation. The work [197] proposes Gaussian Surfels with depth-normal consistency loss and volumetric cutting for improved surface reconstruction, followed by screened Poisson mesh generation. Similarly, 2D Gaussian Splatting [199] (2DGS) replaces 3DGS with planar disks to represent surfaces, which are defined within the local tangent plane. Then, Gaussian Opacity Fields (GOF) [198] are developed based on 3DGS, wherein 3DGS is normalized along the ray to form a 1DGS for volume rendering. Although these methods achieve precise geometry reconstruction, they inevitably compromise rendering fidelity and face challenges in handling semi-transparent surfaces.
V-C Editable 3D Gaussian Splatting